# Pareto Control Barrier Function for Inner Safe Set Maximization Under Input Constraints

### v1

@@ Abstract
This article introduces the Pareto Control Barrier Function (PCBF) algorithm to maximize the inner safe set of dynamical systems under input constraints. Traditional Control Barrier Functions (CBFs) ensure safety by maintaining system trajectories within a safe set but often fail to account for realistic input constraints. To address this problem, we leverage the Pareto multi-task learning framework to balance competing objectives of safety and safe set volume. The PCBF algorithm is applicable to high-dimensional systems and is computationally efficient. We validate its effectiveness through comparison with Hamilton-Jacobi reachability for an inverted pendulum and through simulations on a 12-dimensional quadrotor system. Results show that the PCBF consistently outperforms existing methods, yielding larger safe sets and ensuring safety under input constraints.
@@ I INTRODUCTION
However, ensuring safety becomes challenging when systems operate under input constraints [4, 5, 6, 7]. Traditional CBFs typically assume the availability of arbitrary control inputs [8, 9], which allows the system to be guided within the safe set without limitations. However, this assumption often overlooks the practical reality that control inputs are constrained by physical limitations of actuators, energy resources, or other system-specific constraints. As a result, the safe set may no longer be forward invariant under such restricted inputs [4].
@@
CBFs under input constraints: Recent research has tried to address this problem. Agrawal and Panagou [4] introduced Input Constrained CBFs (ICCBFs), iteratively defining functions to form an inner safe set $C^{*}$ that is forward invariant under input constraints. Breeden and Panagou [5] introduced Zeroing CBFs (ZCBFs) for high-relative-degree systems, ensuring forward invariance by converting higher-order constraints into ZCBFs. However, both approaches tend to be overly conservative, as they do not guarantee the largest possible safe set, potentially limiting the system’s operational flexibility.
@@
Neural CBFs: Neural CBFs (NCBFs) have emerged as an effective approach for ensuring safety in complex, high-dimensional systems, where traditional CBFs may be difficult to design [10]. These methods have been extended to handle parametric uncertainties, unknown dynamic environments, and safe navigation tasks [11, 12, 13]. However, few works have integrated input constraints into NCBFs. Liu et al. proposed the non-saturating CBF (NSCBF) to handle input constraints solving a min-max optimization [6]. Though effective, the complexity of this approach leads to longer training times. Oswin et al. introduced a policy NCBF based on policy iteration [7]. However, their method doesn’t guarantee the largest possible inner safe set either.
@@
HJ reachability: Hamilton-Jacobi (HJ) reachability analysis theoretically computes the largest control-invariant set [14]. Recent work has combined HJ reachability with CBFs to mitigate the overly conservative problem of HJ policies [15]. However, HJ reachability is computationally intensive and practically limited to systems with fewer than five dimensions, restricting its applicability in higher-dimensional dynamical systems.
We summarize our contributions as follows:
@@
  • Maximizing Inner Safe Set: We propose the PCBF algorithm (Algorithm 1) to determine the largest inner safe set under input constraints.
@@
  • Reduced Training Time: We enhance training efficiency by applying a Gaussian sampling method to the inner safe set’s interior.
@@
  • Validation on Benchmark Systems: We validate our method by comparing PCBF results for the inverted pendulum with HJ reachability and through simulations on a 12-dimensional quadrotor system.
@@
The remainder of this article is organized as follows: Section II provides preliminaries, including problem definition, CBFs, Neural CBFs, and Pareto Multi-task Learning. Section III presents our methodology, detailing the PCBF algorithm. Section IV demonstrates the effectiveness of our approach through experiments on an inverted pendulum and a quadrotor system. Finally, Section V concludes the paper and discusses future work directions.
@@ II-A Problem Definition
We consider continuous-time, control-affine dynamics:
@@
where $x\in\mathcal{X}\subseteq\mathbb{R}^{n}$, $u\in\mathcal{U}\subseteq\mathbb{R}^{m}$, and $f,g$ are locally Lipschitz continuous functions. To ensure safety, we introduce the concept of a safe set.
@@
The safe set $S$ is defined to include all states $x$ in which the system remains safe during operation. In practice, the safe set often comes from the physical constraints and safety requirements inherent in the system, such as avoiding collisions or maintaining stability. However, due to the presence of input constraints, ensuring forward invariance of the entire safe set is challenging [4]. Thus, the concept of an inner safe set is introduced.
@@
The goal of this research is to find the largest such inner safe set given the system’s input constraints. The problem is formulated as below:
@@ II-B Control Barrier Functions
CBFs provide a systematic approach to ensuring the forward invariance of safe sets by applying the following lemma [1]:
Lemma 1 (Ames et al. [1]).
Given a set $C=\{x\in\mathcal{X}:h(x)\geq 0\}$ defined by a CBF $h(x)$, the set $C$ is forward invariant under the control input $u\in\mathcal{U}$ if there exists a feedback controller $\pi(x)$ such that:
$L_{f}h(x)+L_{g}h(x)\pi(x)\geq-\alpha(h(x)),\quad\forall x\in C.$ (4)
The CBF Quadratic Program (CBF-QP) is formulated to find the optimal control input $u(x)$ that minimizes the deviation from the desired control $u_{\text{des}}$while satisfying the CBF’s safety constraints:
@@
[…]
However, when input constraints exist, the control input $u$ satisfying (3) may not always comply with the constraints. Certain states within the safe set may no longer be forward invariant. Therefore, the challenge becomes identifying the largest subset of the safe set, the maximum inner safe set.
@@ II-C Neural CBFs
NCBFs address the difficulty of hand-designing CBFs for complex systems by using neural networks to approximate a valid CBF $h_{\theta_{1}}(x)$ and its corresponding controller $\pi_{\theta_{2}}(x)$, where $\theta_{1}$ and $\theta_{2}$ are the parameters of the neural networks for the CBF and controller, respectively [10, 16]. For a given safe set $\mathcal{S}$, NCBFs define three loss functions: $\mathcal{L}_{\text{safe}}$, $\mathcal{L}_{\text{unsafe}}$, and $\mathcal{L}_{\text{feas}}$. To compute these losses, we sample a set of points $\mathbb{X}=\{x_{1},\dots,x_{N}\}\subset\mathcal{X}$, where $N$ is the number of sampled points. Let $N_{\text{safe}}$ and $N_{\text{unsafe}}$ denote the number of sampled points from the safe set $\mathcal{S}$ and the complement of the safe set $\mathcal{X}\setminus\mathcal{S}$.
$\mathcal{L}_{\text{safe}}=\frac{1}{N_{\text{safe}}}\sum_{x_{i}\in\mathbb{X}\cap\mathcal{S}}\max\{0,-h_{\theta_{1}}(x_{i})\},$ (6)
$\mathcal{L}_{\text{unsafe}}=\frac{1}{N_{\text{unsafe}}}\sum_{x_{i}\in\mathbb{X}\cap(\mathcal{X}\setminus\mathcal{S})}\max\{0,h_{\theta_{1}}(x_{i})\},$ (7)
$\mathcal{L}_{\text{feas}}=\frac{1}{N_{\text{safe}}}\sum_{x_{i}\in\mathbb{X}\cap\mathcal{S}}\max\{0,-[L_{f}h_{\theta_{1}}(x_{i})+\\ L_{g}h_{\theta_{1}}(x_{i})\pi_{\theta_{2}}(x_{i})+\alpha(h_{\theta_{1}}(x_{i}))]\}.$ (8)
$\mathcal{L}_{\text{safe}}$ and $\mathcal{L}_{\text{unsafe}}$ penalize states where $h_{\theta_{1}}(x)$ misclassifies safety, either within the safe set or outside it, while $\mathcal{L}_{\text{feas}}$ ensures $\pi_{\theta_{2}}(x)$ satisfies the feasibility condition (4) for forward invariance.
However, such NCBFs still require a predefined safe set. Liu et al. [6] proposed the NSCBF $h_{\theta}(x)$, which removes the need for a predefined safe set and directly handles input constraints:
@@
where $\text{nn}_{\theta}(x)$ is a neural network with parameter $\theta\in\mathbb{R}^{p}$, and $x_{e}$ is a known safe state that must be included in the learned inner safe set. Typically, in problems aimed at maintaining the system’s proximity to a stable point, $x_{e}$ is the stable point itself. For obstacle avoidance problems, $x_{e}$ can be a point far from the obstacle. This method ensures that the learned inner safe set $C_{h_{\theta}}=\{x\in\mathcal{X}:h_{\theta}(x)\geq 0\}$ is always a subset of the original safe set. Also, $C_{h_{\theta}}$ is always a non-empty set, guaranteeing that there are sufficient sample points during the training process.
Ensuring that the learned inner safe set satisfies input constraints, Liu et al. [6] calculate the saturation risk by evaluating the worst-case violation of safety constraints along the boundary of $C_{h_{\theta}}$. To increase the volume of the inner safe set, a regularization term is introduced to encourage the expansion of the safe set. Finally, the total loss is a linear combination of the saturation risk and the regularization term.
Although this method has shown success in specific cases, it has several limitations. First, sampling on $\partial C_{h_{\theta}}$ to calculate the saturation risk is computationally expensive; second, the regularization term does not guarantee finding the largest possible safe set. Even for the maximal inner safe set, the regularization term remains non-zero, which can lead to suboptimal solutions.
@@ III METHODOLOGY
We introduce the Pareto multi-task learning framework [17] into traditional NCBFs to propose the PCBF algorithm. This approach achieves a balance between ensuring safety and maximizing the inner safe set under input constraints. We use the NCBF formulation from equation (9) in this study.
@@ III-A Loss Function Design
$\mathcal{L}_{\text{feas}}=\frac{1}{N_{\mathcal{N}}}\sum_{x_{i}\in\mathbb{X}_{\mathcal{N}}\cap C_{h_{\theta}}}\max\{0,-\sup_{u\in\mathcal{U}}\phi_{\theta}(x_{i},u)\},$ (10)
@@
For simplicity, we assume linear input constraints, making $\mathcal{U}$ a polyhedron. Since $\phi_{\theta}(x,u)$ is affine in $u$, the supremum over $\mathcal{U}$ is necessarily achieved at some vertex of $\mathcal{U}$, allowing closed-form computation of $\mathcal{L}_{\text{feas}}$.
@@
where $\mathcal{V}(\mathcal{U})$ is the set of vertexes of $\mathcal{U}$. The volume loss $\mathcal{L}_{\text{vol}}$ is designed to encourage the expansion of the learned inner safe set by minimizing the deviation from $\mathcal{S}$:
$\mathcal{L}_{\text{vol}}=\frac{1}{N_{\mathcal{N}}}\sum_{x_{i}\in\mathbb{X}_{\mathcal{N}}\cap C_{h_{\theta}}}(\text{nn}_{\theta}(x_{i})-\text{nn}_{\theta}(x_{e}))^{2},$ (13)
However, a simple linear combination of these loss functions does not guarantee a maximized inner safe set. The feasibility loss and the volume loss are inherently competing as both cannot be zero simultaneously—expanding the safe set violates safety constraints, and satisfying constraints reduces the set’s volume.
To address this challenge, the Pareto multi-task learning framework is introduced to balance these conflicting objectives.
@@ III-B Pareto Multi-task Learning (PMTL)
where $\mathcal{L}_{i}(\theta)$ is the loss of the $i$-th task. In general, it is impossible to find a single solution $\theta$ that simultaneously optimizes all the objectives. Instead, we can identify a collection of Pareto optimal solutions, each representing a distinct optimal trade-off among the various objectives [17].
Definition 5 (Pareto Dominance).
@@
Definition 6 (Pareto Optimality).
@@
Definition 7 (Pareto Front).
@@
The Pareto front represents the best possible trade-offs among the competing objectives that can be achieved by any solution. PMTL provides a framework for simultaneously optimizing multiple objectives and converge to the Pareto front. Two common PMTL methods are:
Linear combination: A non-negative linear weighted sum is utilized to aggregate the losses of all tasks into a single loss: $\mathcal{L}(\theta)=\sum_{i=1}^{M}\lambda_{i}\mathcal{L}_{i}(\theta)$, where $\lambda_{i}\geq 0$ is the weight for the $i$-th task. While straightforward, this approach has some limitations: firstly, selecting appropriate values for $\lambda_{i}$ is challenging and often requires extensive empirical tuning; secondly, this method can only yield solutions on the convex portion of the Pareto front [18].
Figure 1: Pareto front of the Inner Safe Set Maximization Problem. The blue line represents the Pareto front, which delineates the boundary between the feasible and infeasible regions in the objective space. No solution $\theta$ can produce objective values $[\mathcal{L}_{\text{feas}}(\theta),\mathcal{L}_{\text{vol}}(\theta)]^{T}$ that lie below this curve. The vertical axis represents solutions with zero feasibility loss. All feasible CBFs lie on this axis. The optimal CBF, with parameter $\theta^{*}$, lies at the intersection of the Pareto front and the vertical axis.
@@
A solution $\theta$ is said to be Pareto critical if there is no solution in its neighborhood can have better performance in all tasks. If $\theta_{t}$ is Pareto critical, $d_{t}=0$ and $\theta_{t}$ will not be updated. If $\theta_{t}$ is not Pareto critical, then $d_{t}$ is a descent directrion for all tasks.
III-C Pareto Control Barrier Function (PCBF)
Problem 1 can be reformulated as a PMTL problem involving two competing objectives: volume loss $\mathcal{L}_{\text{vol}}$ and feasibility loss $\mathcal{L}_{\text{feas}}$. The desired Pareto optimal solution, denoted as $\theta^{*}$, achieves zero feasibility loss while minimizing the volume loss:
@@ III-C Pareto Control Barrier Function (PCBF)
Let $\theta^{*}$ be the solution to problem 19, then $\theta^{*}$ is Pareto optimal, and the loss vector $[\mathcal{L}_{\mathrm{feas}}(\theta^{*}),\mathcal{L}_{\mathrm{vol}}(\theta^{*})]^{T}$ lies on the Pareto front.
Lemma 3 establishes a direct connection between the inner safe set maximization problem 19 and the PMTL framework. Figure 1 illustrates the trade-offs between the two losses on the Pareto front.
The PCBF algorithm uses gradient-based multi-objective optimization methods to efficiently converge to the optimal parameter $\theta^{*}$. It has two main phases: initialization and parameter update.
  • Initialization: The $\text{nn}_{\theta}(\cdot)$ model in equation (9) is trained with a linear combination loss $\mathcal{L}_{\text{lc}}(\theta_{t})=\lambda_{1}\mathcal{L}_{\text{feas}}(\theta_{t})+\lambda_{2}\mathcal{L}_{\text{vol}}(\theta_{t})$ from a randomly initialized $\theta_{0}\in\mathbb{R}^{p}$. Here $\lambda_{1},\lambda_{2}\geq 0$ are hand-designed parameters. At each iteration $t$, parameters are updated by: $\theta_{t+1}=\theta_{t}-\eta\nabla\mathcal{L}_{\text{lc}}(\theta_{t})$. This phase provides an initial solution close to $\theta^{*}$ for further optimization.
  • Parameter update: In this phase, the parameter $\theta_{t}$ is updated iteratively by $\theta_{t+1}=\theta_{t}+\eta d_{t}$. The update direction $d_{t}$ is determined by solving multiple instances of problem (16) with different loss configurations.
Objective space decomposition: As shown in Figure 2, in the parameter update phase, the objective space is partitioned into three regions using two parallel lines: the upper bound $\mathcal{L}_{\text{vol}}(\theta)=\beta\mathcal{L}_{\text{feas}}(\theta)+\epsilon_{\text{ub}}$ and the lower bound $\mathcal{L}_{\text{vol}}(\theta)=\beta\mathcal{L}_{\text{feas}}(\theta)+\epsilon_{\text{lb}}$, where $\beta$, $\epsilon_{\text{ub}}$, and $\epsilon_{\text{lb}}$ are hyperparameters. The key idea of the PCBF is restricting the loss vector between these two bounds. To ensure $\mathbf{L}(\theta^{*})$ lies between these bounds, where $\theta^{*}$ is the solution to Problem 19, we impose  [^2] Although the exact value of $\mathcal{L}_{\text{vol}}(\theta^{*})$ is unknown, we can obtain an estimate by first performing a pre-training process using the linear combination loss. Based on this estimate, we can then adjust $\epsilon_{\text{ub}}$ and $\epsilon_{\text{lb}}$ accordingly. the constraints $\beta>0$ and $0\leq\epsilon_{\text{lb}}\leq\mathcal{L}_{\text{vol}}(\theta^{*})\leq\epsilon_{\text{ub}}$. For each region, we solve different subproblems to determine the update direction $d_{t}$:
Figure 2: The blue solid curve represents the Pareto front, while the black dashed lines indicate the upper and lower bounds. The green point denotes a feasible solution (i.e., a solution $\theta$ with $\mathbf{L}(\theta)$ lies between the two bounds), the red point represents an infeasible solution, and the blue point corresponds to $\theta^{*}$, the solution of problem 19. If the update causes $\theta$ to fall outside the feasible region, the PCBF algorithm will gradually guide $\theta$ back into the feasible region bounded by the upper and lower bounds.

### v2

@@ Abstract
Control Barrier Functions (CBFs) enforce safety in dynamical systems by ensuring trajectories stay within prescribed safe sets. However, traditional CBFs often overlook realistic input constraints, potentially limiting their practical effectiveness. To address this research gap, this paper introduces the Pareto Control Barrier Function (PCBF) algorithm, which employs a Pareto multi-task learning framework to simultaneously balance the competing objectives of maintaining safety and maximizing the volume of the safe set under input limitations. The PCBF algorithm is computationally efficient and scalable to high-dimensional systems. The effectiveness of PCBF is demonstrated through comparisons with Hamilton-Jacobi reachability on an inverted pendulum and extensive simulations on a 12-dimensional quadrotor system, where PCBF consistently outperforms existing methods by yielding larger safe sets while ensuring robust safety under input constraints. Codes are available at https://github.com/XiaoyangCao1113/Pareto_CBF.
@@ I INTRODUCTION
However, ensuring safety becomes challenging when systems operate under input constraints [4, 5, 6, 7]. Traditional CBFs typically assume the availability of arbitrary control inputs [8, 9], which allows the system to be guided within the safe set without limitations. Yet, this assumption often ignores the fact that control inputs are constrained by actuator limits, energy resources, and other system-specific factors. As a result, the safe set may no longer be forward invariant under such restricted inputs [4].
@@
Recent research has tried to address this problem. Agrawal and Panagou [4] introduced Input Constrained CBFs (ICCBFs), iteratively defining functions to form an inner safe set $C^{*}$ that is forward invariant under input constraints. Breeden and Panagou [5] introduced Zeroing CBFs (ZCBFs) for high-relative-degree systems, ensuring forward invariance by converting higher-order constraints into ZCBFs. However, both approaches tend to be overly conservative, as they do not guarantee the largest possible safe set, potentially limiting the system’s operational flexibility.
@@
Neural CBFs (NCBFs) have emerged as an effective approach to ensure safety in complex, high-dimensional systems, where traditional CBFs may be difficult to design [10]. These methods have been extended to handle parametric uncertainties, unknown dynamic environments, and safe navigation tasks [11, 12, 13]. However, few works have integrated input constraints into NCBFs. Liu et al. proposed the non-saturating CBF (NSCBF) to handle input constraints solving a min-max optimization [6]. Though effective, the complexity of this approach leads to longer training times. Oswin et al. introduced a policy NCBF based on policy iteration [7]. However, their method does not guarantee the maximal inner safe set either.
@@
Hamilton-Jacobi (HJ) reachability analysis theoretically computes the largest control-invariant set [14]. Recent work has combined HJ reachability with CBFs to mitigate the overly conservative problem of HJ policies [15]. However, HJ reachability is computationally intensive and practically limited to systems with fewer than five dimensions, restricting its applicability to higher-dimensional dynamical systems.
To address these limitations, we make the following contributions:
@@
  • We propose the PCBF algorithm to determine the largest inner safe set under input constraints.
@@
  • We enhance training efficiency by applying a Gaussian sampling method to the inner safe set’s interior.
@@
  • We validate the proposed approach through comparisons with HJ reachability on an inverted pendulum and demonstrate its scalability via simulations on a 12-dimensional quadrotor system.
@@
The remainder of this article is organized as follows: Section II provides preliminaries, including problem definition, CBFs, Neural CBFs, and Pareto Multi-task Learning. Section III presents our methodology, detailing the PCBF algorithm. Section IV demonstrates the effectiveness of our approach through experiments on an inverted pendulum and a quadrotor system. Section V concludes the paper and discusses future work directions.
@@ II PRELIMINARIES
In this section, we present the theoretical framework of CBFs and Neural CBFs, and how they ensure safety guarantees.
@@ II-A Problem Definition
Consider continuous-time, control-affine dynamics:
@@
where $x\in\mathcal{X}\subseteq\mathbb{R}^{n}$, $u\in\mathcal{U}\subseteq\mathbb{R}^{m}$, and $f,g$ are locally Lipschitz continuous functions. To ensure safety, the concept of a safe set is introduced.
@@
The safe set $S$ is defined to include all states $x$ in which the system remains safe during operation. In practice, the safe set often comes from the physical constraints and safety requirements inherent in the system, such as avoiding collisions or maintaining stability. However, the presence of input constraints makes it challenging to ensure forward invariance of the entire safe set [4]. Thus, the concept of an inner safe set is introduced.
@@
This research aims to find the largest such inner safe set given the system’s input constraints. The problem is formulated as below:
@@ II-B Control Barrier Functions
Control Barrier Functions (CBFs) provide a systematic approach to ensuring the forward invariance of safe sets.
@@
To compute a safe control input $u(x)$, the following CBF Quadratic Program (CBF-QP) [16] is formulated. The objective is to find a control input that remains as close as possible to the nominal control policy $\pi_{nom}(x):\mathcal{X}\rightarrow\mathcal{U}$ while enforcing the safety constraints. According to Theorem 2 in [1], the solution to problem 4 guarantees the forward invariance of the safe set.
@@
The nominal policy $\pi_{nom}$ is derived using unconstrained control design methods (e.g., LQR, MPC, etc.) that do not inherently guarantee safety. For instance, consider a self-driving car following a planned route. Its nominal policy is the ideal steering and acceleration commands for smooth lane tracking. However, if a pedestrian suddenly steps onto the road, the CBF-QP adjusts the commands from $\pi_{nom}$ just enough to safely avoid a collision, while still trying to stay close to the original plan.
When $\mathcal{U}$ is unbounded, the affine structure of (3) in $u$ ensures that a valid CBF and a corresponding control input $u$ can always be found, provided $L_{g}h(x)\not\equiv 0$. However, under input constraints, the control input $u$ that satisfies (3) may no longer remain within the admissible set $\mathcal{U}$. As a result, certain states within the safe set may lose forward invariance. This raises the key challenge of identifying the largest inner safe set that is compatible with input constraints.
@@ II-C Neural CBFs
Neural CBFs (NCBFs) address the difficulty of hand-designing CBFs for complex systems by using neural networks to approximate a CBF and its corresponding controller [10, 16]. However, such NCBFs still require a predefined safe set. If this safe set contains any unsafe states, the neural networks will fail to converge. Liu et al. [6] proposed the NSCBF $h_{\theta}(x)$, which removes the need for a predefined safe set and incorporates input constraints into its formulation.
@@
where $\text{nn}_{\theta}(x)$ is a neural network with parameter $\theta\in\mathbb{R}^{p}$, and $x_{e}$ is a known safe state. Typically, in problems aimed at maintaining the system’s proximity to a stable point, $x_{e}$ itself is the stable point under some appropriate control law. For obstacle avoidance problems, $x_{e}$ can be a point far from the obstacle. This method ensures that the learned inner safe set $C_{h_{\theta}}=\{x\in\mathcal{X}:h_{\theta}(x)\geq 0\}$ is always a subset of the original safe set. Also, $C_{h_{\theta}}$ is always a non-empty set, guaranteeing that there are sufficient sample points during the training process.
Although this method has shown success in specific cases, it has several limitations. First, it requires sampling on $\partial C_{h_{\theta}}$ to calculate the saturation risk, which is computationally expensive; second, it does not guarantee the identification of the largest possible safe set. Moreover, even for the maximal inner safe set, the loss function remains non-zero, which can lead to suboptimal solutions.
@@ III METHODOLOGY
The Pareto multi-task learning framework [17] is introduced into traditional NCBFs to develop the PCBF algorithm. This approach maintains a balance between ensuring safety and maximizing the inner safe set under input constraints. The NCBF formulation from equation (5) is adopted throughout this study.
@@ III-A Loss Function Design
Feasibility loss: The feasibility loss $\mathcal{L}_{\text{feas}}$ is necessary for the forward invariance of the learned inner safe set $C_{h_{\theta}}$:
$\mathcal{L}_{\text{feas}}=\frac{1}{N_{\theta}}\sum_{x_{i}\in\mathbb{X}_{\mathcal{N}}\cap C_{h_{\theta}}}\max\{0,-\sup_{u\in\mathcal{U}}\phi_{\theta}(x_{i},u)\},$ (6)
@@
where $N_{\theta}$ is the number of sampling points in $\mathbb{X}_{\mathcal{N}}\cap C_{h_{\theta}}$. For simplicity, this study assumes linear input constraints, making $\mathcal{U}$ a polyhedron. Since $\phi_{\theta}(x,u)$ is affine in $u$, the supremum over $\mathcal{U}$ is necessarily achieved at some vertex of $\mathcal{U}$, allowing closed-form computation of $\mathcal{L}_{\text{feas}}$.
@@
where $\mathcal{V}(\mathcal{U})$ is the set of vertexes of $\mathcal{U}$.
Volume loss: The volume loss $\mathcal{L}_{\text{vol}}$ is designed to encourage the expansion of the learned inner safe set by minimizing the deviation from $\mathcal{S}$:
$\mathcal{L}_{\text{vol}}=\frac{1}{N_{\theta}}\sum_{x_{i}\in\mathbb{X}_{\mathcal{N}}\cap C_{h_{\theta}}}(\text{nn}_{\theta}(x_{i})-\text{nn}_{\theta}(x_{e}))^{2},$ (9)
However, a simple linear combination of these loss functions does not guarantee a maximized inner safe set. The feasibility loss and the volume loss are inherently competing as both cannot be zero simultaneously: Expanding the safe set violates safety constraints, and satisfying constraints reduces the set volume. To address this challenge, the Pareto multi-task learning framework is introduced to balance these conflicting objectives.
Figure 1: Pareto front of the Inner Safe Set Maximization Problem. The blue line represents the Pareto front, which delineates the boundary between the feasible and infeasible regions in the objective space. No solution $\theta$ can produce objective values $[\mathcal{L}_{\text{feas}}(\theta)\quad\mathcal{L}_{\text{vol}}(\theta)]^{T}$ that lie below this curve. The vertical axis represents solutions with zero feasibility loss. All feasible CBFs lie on this axis. The optimal CBF, with parameter $\theta^{*}$, lies at the intersection of the Pareto front and the vertical axis.
@@ III-B Pareto Multi-task Learning (PMTL)
where $\mathcal{L}_{i}(\theta)$ is the loss of the $i$-th task. In general, it is impossible to find a single solution $\theta$ that simultaneously optimizes all the objectives. Instead, a collection of Pareto optimal solutions is identified, each representing a distinct optimal trade-off among the various objectives [17].
Definition 5 (Pareto Dominance [18]).
@@
Definition 6 (Pareto Optimality [18]).
@@
Definition 7 (Pareto Front [18]).
@@
The Pareto front represents the set of optimal trade-offs among competing objectives achievable by any solution. PMTL provides a framework for simultaneously optimizing multiple objectives and converges to the Pareto front. Two common PMTL methods are:
Linear combination: A non-negative linear weighted sum is utilized to aggregate the losses of all tasks into a single loss: $\mathcal{L}(\theta)=\sum_{i=1}^{M}\lambda_{i}\mathcal{L}_{i}(\theta)$, where $\lambda_{i}\geq 0$ is the weight for the $i$-th task. Although straightforward, this approach has some limitations: first, selecting appropriate values for $\lambda_{i}$ is challenging and often requires extensive empirical tuning; second, this method can only produce solutions on the convex portion of the Pareto front [19].
@@
Figure 2: The blue solid curve represents the Pareto front, while the black dashed lines indicate the upper and lower bounds. The green point denotes a feasible solution (i.e., a solution $\theta$ with $\mathbf{L}(\theta)$ lies between the two bounds), the red point represents an infeasible solution, and the blue point corresponds to $\theta^{*}$, the solution of problem 15. If the update causes $\theta$ to fall outside the feasible region, the PCBF algorithm will gradually guide $\theta$ back into the feasible region bounded by the upper and lower bounds.
@@
A solution $\theta$ is said to be Pareto critical if no solution in its neighborhood can perform better in all tasks. If $\theta_{t}$ is Pareto critical, $d_{t}=0$ and $\theta_{t}$ will not be updated. If $\theta_{t}$ is not Pareto critical, then $d_{t}$ is a descent direction for all tasks.
Connection to problem 1: Problem 1 can be reformulated as a PMTL problem with $M=2$ and $\mathbf{L}(\theta)=[\mathcal{L}_{\mathrm{feas}}(\theta)\quad\mathcal{L}_{\mathrm{vol}}(\theta)]^{T}$ emperically.
@@
The solution of problem 15, denoted as $\theta^{*}$, achieves zero feasibility loss while minimizing the volume loss, and lemma 2 shows that $\theta^{*}$ is on the Pareto front. Figure 1 provides a visual explanation of the trade-offs between the two losses on the Pareto front.
@@
Let $\theta^{*}$ be the solution to problem 15, then $\theta^{*}$ is Pareto optimal, and the loss vector $[\mathcal{L}_{\mathrm{feas}}(\theta^{*})\mathcal{L}_{\mathrm{vol}}(\theta^{*})]^{T}$ lies on the Pareto front.
The PCBF algorithm is proposed to effectively solve problem 15 by applying gradient-based PMTL methods.
III-C Pareto Control Barrier Function (PCBF)
Initially, we attempted to employ a linear combination of $\mathcal{L}_{\text{feas}}$ and $\mathcal{L}_{\text{vol}}$ as the total loss function. However, this approach often led to one of two outcomes: either $\mathcal{L}_{\text{vol}}\approx 0$ (resulting in a model that fails to remove potentially hazardous states) or $\mathcal{L}_{\text{feas}}\approx 0$ but with a large $\mathcal{L}_{\text{vol}}$ (yielding an excessively small inner safe set).
To mitigate these issues, the objective space is partitioned into three regions using two parallel lines (as shown in Figure 2). The intermediate region is designed to contain $\mathbf{L}(\theta^{*})$. The motivation behind the PCBF algorithm is to restrict $\mathbf{L}(\theta)$ within this intermediate region, thereby avoiding convergence to the aforementioned extreme outcomes while guiding the gradient descent towards $\theta^{*}$. Specifically, we define an upper bound: $\mathcal{L}_{\text{vol}}(\theta)=\beta\mathcal{L}_{\text{feas}}(\theta)+\epsilon_{\text{ub}},$ and a lower bound: $\mathcal{L}_{\text{vol}}(\theta)=\beta\mathcal{L}_{\text{feas}}(\theta)+\epsilon_{\text{lb}},$ where $\beta$, $\epsilon_{\text{ub}}$, and $\epsilon_{\text{lb}}$ are hyperparameters. To ensure $\mathbf{L}(\theta^{*})$ lies between these bounds, constraints are imposed  [^2] Although the exact value of $\mathcal{L}_{\text{vol}}(\theta^{*})$ is unknown, we can obtain an estimate by first performing a pre-training process using the linear combination loss. such that $\beta>0$ and $0\leq\epsilon_{\text{lb}}\leq\mathcal{L}_{\text{vol}}(\theta^{*})\leq\epsilon_{\text{ub}}$.
In iteration $t+1$, the parameter update is given by: $\theta_{t+1}=\theta_{t}+\eta d_{t}$, where $d_{t}$ is the update direction and $\eta>0$ is the learning rate. To ensure that the updated parameters remain within the specified region, $d_{t}$ is determined based on the position of $\mathbf{L}(\theta_{t})$ in the objective space: